OpenStudy (anonymous):

Given a+b+c=0, prove that...

OpenStudy (anonymous):

\[a(\frac{ 1 }{ b }+\frac{ 1 }{ c })+b(\frac{ 1 }{ c }+\frac{ 1 }{ a })+c(\frac{ 1 }{ a }+\frac{ 1 }{ b })+3=0\]